# Exploiting Euclidean Distance Field Properties for Fast and Safe 3D planning with a modified Lazy Theta*

### v1

@@ Abstract
Graph search planners have been widely used for 3D path planning in the literature, and Euclidean Distance Fields (EDFs) are increasingly being used as a representation of the environment. However, to the best of our knowledge, the integration of EDFs into heuristic planning has been carried out in a loosely coupled fashion, dismissing EDF properties that can be used to accelerate/improve the planning process and enhance the safety margins of the resultant trajectories. This paper presents a fast graph search planner based on a modified Lazy Theta* planning algorithm for aerial robots in challenging 3D environments that exploits the EDF properties. The proposed planner outperforms classic graph search planners in terms of path smoothness and safety. It integrates EDFs as environment representation and directly generates fast and smooth paths avoiding the use of post-processing methods; it also considers the analytical properties of EDFs to obtain an approximation of the EDF cost along the line-of-sight segments and to reduce the number of visibility neighbours, which directly impacts the computation time. Moreover, we demonstrate that the proposed EDF-based cost function satisfies the triangle inequality, which reduces calculations during exploration and, hence, computation time. Many experiments and comparatives are carried out in 3D challenging indoor and outdoor simulation environments to evaluate and validate the proposed planner. The results show an efficient and safe planner in these environments.
@@ 1.1 Motivation
Path planning in complex 3D environments is a basic procedure that robots must carry out frequently in numerous nowadays applications: tunnel inspection [1], urban firefighting [2], subterranean inspection, like the DARPA Sub-T challenge111 https://www.darpa.mil/program/darpa-subterranean-challenge or drone racing challenges such as AlphaPilot [3]. Approaches ranging from sampling-based to searching-based planners have been successfully applied in the state of the art. We can also find methods based on curve interpolation and optimization. Graph search and heuristic planners such as A* [4] or Theta* [5] are widely used. These methods usually focus on the generation of optimal paths with respect to the path length, leaving aside safety and smoothness of the path (see Fig. 1). As a consequence, computed paths tend to include unnecessary heading changes, which requires post-processing methods [6, 7] to mitigate such effect.
@@
Euclidean Distance Fields (EDF) are a popular environment representation. The field gives the distance to the closest obstacle for any point in space. EDFs have interesting analytical properties [8]. They are particularly useful in 3D path planning [9, 10, 11, 12], but to the best for our knowledge, most existing path planning approaches that integrate EDFs do not exploit their analytical properties. Despite the usefulness of EDF in path planning, heuristic search planners are rarely integrated with them to leverage their properties [13]. Fig. 1 shows the different paths computed by the A* algorithm using a grid map as obstacle representation and the path length as cost, and the A*+M1 [13] which considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. Notice how considering distance to obstacles using EDF as representation enables safer and smoother paths [13]. On the other hand, the path length as cost holds the triangle inequality, which states that the sum of the lengths of two sides of a triangle must be longer than the length of the remaining side. It can be used in heuristic planners to simplify the calculations to select the parent node.
@@
Figure 1: Paths computed by original A* algorithm in blue and red, and computed by A*+M1 from [13] in blue considering distance to obstacles represented as EDF: 2D (left) and 3D (right). Both paths show how the distance to obstacles influences on the computed path by A*+M1.
@@
The goal of this paper is to develop efficient and safe heuristic path planners able to integrate an EDF, consider an EDF-based cost function and take advantage of its properties in order to directly compute safe and smooth paths. It will be also shown that we can even modify existing heuristic planners as both A* and Lazy Theta* planning algorithms to naturally exploit the EDF properties. This will enable approximation of the EDF cost along a segment and reduce the number of visibility neighbours, which affects the computation time. Furthermore, post-processing methods are no longer required to smooth the path. Finally, if we can demonstrate that the proposed EDF-based cost function holds the triangle inequality, the simplification of the calculations will also be ensured. This would be an extension of the previous work [13] because its cost function does not satisfy the triangle inequality.
@@ 1.2 Related work
Oleynikova et al. [9] use Euclidean Signed Distance Field (ESDF) to consider the distance cost to the obstacles using Informed RRT*, but the planned trajectory can vary significantly from execution to execution and the EDF properties are not exploited. The authors in [19] adopt a fast marching path planning method on velocity fields using ESDF, but they use Bernstein basis polynomials to generate smooth paths. Moreover, fast marching methods tend to find longer paths than searching-based algorithms as A* since the metric is velocity in the field, not the path length. EDFs are also integrated as environment representation in heuristic planners in the previous work of the authors [13] so that smooth and safe paths are directly generated with a Lazy Theta* algorithm.
@@ 1.3 Contributions
    A modified Lazy Theta* algorithm with a novel exploration of the visibility neighbours that foster velocity and reduce the need of post-processing methods.
This paper extends the authors’ previous work [13] by further leveraging the EDF properties, which lead to modifications in the Lazy Theta* algorithm. Additionally, a new cost function is introduced to ensure that the triangle inequality holds. A fully detailed demonstration to approximate the cost along the line-of-sight segments on EDFs is presented, and a new modified Lazy Theta* algorithm is proposed. The exploration of visibility neighbours is significantly reduced by considering the obstacle-gradient information, producing faster, smoother and safer paths than those obtained with [13]. How visibility neighbours are selected also affects safety and smoothness.
The planner is evaluated with many random tests in challenging 3D environments. The implementation of the algorithm and the scripts to launch the tests are open sourced222 https://github.com/robotics-upo/Heuristic_path_planners/.
@@ 2 Fast and Safe Path Planner Overview
The proposed planner integrates an EDF to efficiently represent the robot’s environment and to define the graph employed in the planner. The 3D EDF voxel grid is generated offline considering the distance to the closest occupied node for every node in a map with fixed resolution. As we will see, due to the EDF properties, we can obtain an approximation of the EDF cost along segments (see Section 3) and the EDF representation enables easy obstacle gradient computation based on local EDF derivative (see Section 4).
Algorithm 1 shows the modifications performed in the original Lazy Theta* algorithm (highlighted in red). First of all, a cost function that takes into account the distance to obstacles from the EDF (see $ComputeCost$ function of Algorithm 1) is used with the aim of computing safer paths (the cost function is described in Section 3). Another necessary modification is to limit the line of sight ($LoS$) in the $SetVertex$ function of Algorithm 1 in order to generate smoother paths and foster safety. Otherwise, different nodes of the path could be linked if they are visible to each other, overriding the possible benefits of introducing the distance cost term in order to generate safer paths. These modifications imply an extension of the algorithm rather than a modification on how it works.
1 Main()
@@
4       parent($s_{start}$):=$s_{start}$;
5       open.Insert($s_{start}$,$g(s_{start})+h(s_{start})$);
@@ 45             If $s^{\prime}\in$ open then
49            open.Insert($s^{\prime}$,$g(s^{\prime})+h(s^{\prime})$);
@@
56       If $g(parent(s))+||s^{\prime}-parent(s)||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}}<g(s^{\prime})$ then
@@ 57             $parent(s^{\prime})$ := $parent(s)$;
58             $g(s^{\prime})$ := $g(parent(s))+||s^{\prime}-parent(s)||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}}$ ;
@@
64       If NOT lineofsight(parent(s),s) OR lineofsight(parent(s),s) $>$ max then
@@
66             $parent(s)$ := $\arg\min_{s^{\prime}\in nghbr_{vis}\cap closed}(g(s^{\prime})+||s^{\prime}-s||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}})$;
67             $g(s)$ := $\min_{s^{\prime}\in nghbr_{vis}\cap closed}(g(s^{\prime})+||s^{\prime}-s||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}})$;
@@ 3 The cost function
$c(s_{i},s_{i+1})=||s_{i+1}-s_{i}||+\frac{c_{w}}{O(s_{i},s_{i+1})}$ (2)
where $||s_{i+1}-s_{i}||$ is the Euclidean distance between $s_{i}$ and $s_{i+1}$ (the usual term considered in the original Lazy Theta*), $c_{w}$ is a cost weight factor and $O(s_{i},s_{i+1})$ is the integral of the EDF along the segment defined by $s_{i}$ and $s_{i+1}$.
@@
3.1 Computation of $O(s_{i},s^{\prime}_{i+1})$
@@
An important property of the distance function is that it is Lipschitz continuous with Lipschitz constant equal to 1 [21, Proposition 1.74], and its gradient norm is $\|\nabla_{x}d(x;\Omega)\|=1$.
@@
$\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||-\frac{||s_{i+1}-s_{i}||^{2}}{2}\leq O(s_{i},s^{\prime}_{i+1})\leq\\ \leq\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||$ (7)
Proof: First of all, the distance function in $\mathbb{R}^{n}$ with respect to a convex set is a convex function (see [21, Proposition 1.77]). Therefore, $d(l(t);\Omega)$ is a convex function. The Hermite-Hadamard inequality [22] states that if a function $f:[a,b]\rightarrow\mathbb{R}$ is convex the following inequality holds:
@@
$O(s_{i},s^{\prime}_{i+1})\approx\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||$ (11)
As mentioned, the bounds for the distance function apply to obstacles that can be represented by a convex set. In general, a scenario will, however, consist of non-convex obstacles. But it is known that a non-convex set can be represented in general as the union of convex sets [23, 24], and each of them will define a Voronoi region in which the distance function $d(l(t);\Omega)$ is convex. Therefore, the approximation in (11) will be valid except for the segments that cross the boundaries between these regions, the cut locus. Therefore, the maximum allowed length of the segments is an important parameter in the method, as longer lengths will lead to potentially larger errors in the approximation, as the probability of crossing the cut locus increases.
Furthermore, note that the difference between the bounds of (7) is $||s_{i+1}-s_{i}||^{2}/2$ or $L^{2}/2$. This demonstration is relevant because the value of $L$ depends on the line of sight considered, that is, the distance between $s_{i}$ and $s_{i+1}$ nodes will always be less than or equal to the maximum line of sight considered. So, the shorter the line of sight or length of the segments, the more similar the bounds are. This will be evaluated in Section 5.
@@ 4 Selection of visibility neighbours
where $EDF(s)$ and $EDF(s^{\prime})$ are the distances to the closest obstacles in each node $s$ and $s^{\prime}$, and $distance$ is the distance between $s$ and $s^{\prime}$. Notice how positive values indicate that we are getting closer to an obstacle in such direction, while negative ones imply moving away from the obstacle.
Then, the minimum value of $EDF^{\prime}(s^{\prime})$ of the twenty six neighbours, $s_{edf}$, is chosen with the $ComputeMinDerivEDF$ function in Algorithm 1 (see red arrow in Figure 4(b)). On the other hand, the direction from each $s^{\prime}$ to the goal is also computed. The direction of the node $s^{\prime}$ with minimum value of $EDF^{\prime}$ to the goal, $s_{goal}$, is the blue arrow in Figure 4(b). The unit vectors of both $s_{edf}$ and $s_{goal}$ give rise to the direction that defines the best candidate node to expand, $\hat{u}$, thanks to the $ComputeVector$ function in line 16 of Algorithm 1 (black arrow is the direction and light gray box the best candidate to expand in Figure 4(b)). Once $\hat{u}$ is computed, the neighbours are selected around the best candidate with the $ChooseNeighbours$ function in line 17 in Algorithm 1 (gray boxes in Figure 4(b)). For example, Figure 4(b) shows nine neighbours selected instead of twenty six neighbours selected in the original algorithm. Finally, the cost, $g$, and parent of each node is computed for the selected neighbours in line 18-23 in Algorithm 1.
@@
Figure 4: Selection of the visibility neighbours considering the obstacle-gradient information and the goal. The black box represents an obstacle.

### v2

@@ Abstract
This paper presents the FS-Planner, a fast graph-search planner based on a modified Lazy Theta* algorithm that exploits the analytical properties of Euclidean Distance Fields (EDFs). We introduce a new cost function that integrates an EDF-based term proven to satisfy the triangle inequality, enabling efficient parent selection and reducing computation time while generating safe paths with smaller heading variations. We also derive an analytic approximation of the EDF integral along a segment and analyze the influence of the line-of-sight limit on the approximation error, motivating the use of a bounded visibility range. Furthermore, we propose a gradient-based neighbour-selection mechanism that decreases the number of explored nodes and improves computational performance without degrading safety or path quality. The FS-Planner produces safe paths with small heading changes without requiring the use of post-processing methods. Extensive experiments and comparisons in challenging 3D indoor simulation environments, complemented by tests in real-world outdoor environments, are used to evaluate and validate the FS-Planner. The results show consistent improvements in computation time, exploration efficiency, safety, and smoothness in a geometric sense compared with baseline heuristic planners, while maintaining sub-optimality within acceptable bounds. Finally, the proposed EDF-based cost formulation is orthogonal to the underlying search method and can be incorporated into other planning paradigms.
@@ 1.1 Motivation
Path planning in complex 3D environments is a basic procedure that robots must carry out frequently in numerous nowadays applications: tunnel inspection [5], urban firefighting [15], subterranean inspection, like the DARPA Sub-T challenge [^1] https://www.darpa.mil/program/darpa-subterranean-challenge or drone racing challenges such as AlphaPilot [7]. Approaches ranging from sampling-based to searching-based planners have been successfully applied in the state of the art. We can also find methods based on curve interpolation and optimization. Graph search and heuristic planners such as A* [10] or Theta* [1] are widely used. These methods usually focus on the generation of optimal paths with respect to the path length, leaving aside safety and smoothness of the path (see Fig. 1). As a consequence, computed paths tend to include unnecessary heading changes, which requires post-processing methods [6, 14] to mitigate such effect.
@@
Euclidean Distance Fields (EDF) are a popular environment representation. The field gives the distance to the closest obstacle for any point in space. EDFs have interesting analytical properties [13]. They are particularly useful in 3D path planning [20, 8, 22, 25], but to the best for our knowledge, most existing path planning approaches that integrate EDFs do not exploit their analytical properties. Despite the usefulness of EDF in path planning, heuristic search planners are rarely integrated with them to leverage their properties [2]. Fig. 1 shows the different paths computed by the A* algorithm using a grid map as obstacle representation and the path length as cost, and the A*+M1 [2] which considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. Notice how considering distance to obstacles using EDF as representation enables safer and geometrically smoother paths (in terms of heading changes) [2]. On the other hand, the path length as cost holds the triangle inequality, which states that the sum of the lengths of two sides of a triangle must be longer than the length of the remaining side. It can be used in heuristic planners to simplify the calculations to select the parent node.
@@
Figure 1: Paths computed by original A* algorithm in black, and computed by A*+M1 from [2] in purple considering distance to obstacles represented as EDF: 2D (left) and 3D (right). Both paths show how the distance to obstacles influences on the computed path by A*+M1.
@@
The goal of this paper is to develop efficient and safe heuristic path planners able to integrate an EDF, consider an EDF-based cost function and take advantage of its properties in order to directly compute safe and geometrically smooth paths. It will be also shown that we can even modify existing heuristic planners as both A* and Lazy Theta* planning algorithms to naturally exploit the EDF properties. This will enable approximation of the EDF cost along a segment, in addition to using the EDF gradient to identify directions that naturally lead the planner away from obstacles, thereby reducing the number of visibility neighbours and the corresponding computation time. Furthermore, this reduction in neighbour exploration can improve path smoothness with fewer abrupt heading changes, decreasing the need for any post-processing methods. Finally, if we can demonstrate that the proposed EDF-based cost function holds the triangle inequality, which preserves the parent-selection efficiency of Lazy Theta*, the simplification of the calculations will also be ensured. This would be an extension of the previous work [2] because its cost function does not satisfy the triangle inequality and therefore requires additional parent comparisons.
@@ 1.2 Related work
Oleynikova et al. [20] use Euclidean Signed Distance Field (ESDF) to consider the distance cost to the obstacles using Informed RRT*, but the planned trajectory can vary significantly from execution to execution. On the other hand, the proposed path planner is deterministic, producing the same solution under identical conditions. Moreover, [20] uses the EDF solely as a collision-avoidance metric, while the proposed path planner integrates an EDF-based cost function and exploits the analytical properties of the EDF for both cost evaluation and neighbour selection. The authors in [9] adopt a fast marching path planning method on velocity fields using ESDF, but they use Bernstein basis polynomials to generate smooth paths. Moreover, fast marching methods tend to find longer paths than searching-based algorithms as A* since the metric is velocity in the field, not the path length, whereas the proposed path planner directly balances distance, safety, and geometric smoothness. Furthermore, [9] requires a post-processing smoothing stage, while the proposed path planner can generate smooth paths directly. Finally, fast marching methods scale poorly in cluttered 3D environments. EDFs are also integrated as environment representation in heuristic planners in the previous work of the authors [2] so that safe paths with fewer abrupt heading changes are directly generated with a Lazy Theta* algorithm. The advantages of the proposed path planner with respect to [2] are highlighted in Section 1.3.
@@ 1.3 Contributions
    A modified Lazy Theta* algorithm with a novel exploration of the visibility neighbours that foster velocity and improve path smoothness in a geometric sense, which reduces the need for post-processing methods.
This paper extends the authors’ previous work [2] by further leveraging the EDF properties, which lead to modifications in the Lazy Theta* algorithm. First, a distance EDF-based cost function is introduced and satisfies the triangle inequality; in contrast, [2] does not satisfy this property. Second, the proposed path planner reduces the visibility-neighbour set using the EDF gradient, which substantially decreases the number of explored nodes and leads to faster computation, producing geometrically smoother and safer paths than those obtained with [2]. How visibility neighbours are selected also affects safety and smoothness. Additionally, unlike [2], the proposed path planner provides a more detailed theoretical justification of the EDF properties, explicitly covering the case of non-convex obstacles through convex decomposition (see Section 3.1).
The planner is evaluated with many random tests in challenging 3D environments. The implementation of the algorithm and the scripts to launch the tests are open sourced [^2] https://github.com/robotics-upo/Heuristic_path_planners/.
@@ 2 Fast and Safe Path Planner Overview
The proposed planner integrates an EDF to efficiently represent the robot’s environment and to define the graph employed in the planner. The 3D EDF voxel grid is generated offline considering the distance to the closest obstacle for every node in a map with fixed resolution. As we will see, due to the EDF properties, we can obtain an approximation of the EDF cost along segments (see Section 3) and the EDF representation enables easy obstacle gradient computation based on local EDF derivative (see Section 4).
Algorithm 1 shows the modifications performed in the original Lazy Theta* algorithm (highlighted in red). First of all, a cost function that takes into account the distance to obstacles from the EDF (see $ComputeCost$ function of Algorithm 1) is used with the aim of computing safer paths (the cost function is described in Section 3). Another necessary modification is to limit the line of sight ($LoS$) in the $SetVertex$ function of Algorithm 1 in order to generate geometrically smoother paths and foster safety. Otherwise, different nodes of the path could be linked if they are visible to each other, overriding the possible benefits of introducing the distance cost term in order to generate safer paths. These modifications imply an extension of the algorithm rather than a modification on how it works.
1Main()
@@
5  open.Insert($s_{start}$$g(s_{start})+h(s_{start})$);
6  while open $\neq\emptyset$ do
@@ 45     If $s^{\prime}\in$ open then
49    open.Insert($s^{\prime}$$g(s^{\prime})+h(s^{\prime})$);
@@
[…]
@@ 57     $parent(s^{\prime})$ := $parent(s)$;
[…]
@@
64  If NOT lineofsight(parent(s),s) OR lineofsight(parent(s),s) >> max then
@@
[…]
@@ 3 The cost function
[…]
where $s_{i+1}-s_{i}||$ is the Euclidean distance between $s_{i}$ and $s_{i+1}$ (the usual term considered in the original Lazy Theta*), $c_{w}$ is a cost weight factor and $O(s_{i},s_{i+1})$ is the integral of the EDF along the segment defined by $s_{i}$ and $s_{i+1}$.
@@
3.1 Computation of the integral of the EDF along the segment
@@
An important property of the distance function is that it is Lipschitz continuous with Lipschitz constant equal to 1 [16, Proposition 1.74], and its gradient norm is $\|\nabla_{x}d(x;\Omega)\|=1$.
@@
Proof: First of all, the distance function in $\mathbb{R}^{n}$ with respect to a convex set is a convex function (see [16, Proposition 1.77]). Therefore, $d(l(t);\Omega)$ is a convex function. The Hermite-Hadamard inequality [4] states that if a function $f:[a,b]\rightarrow\mathbb{R}$ is convex the following inequality holds:
$(b-a)f(\frac{a+b}{2})\leq\int_{a}^{b}f(x)dx\leq(b-a)\frac{f(a)+f(b)}{2}$ (8)
@@
[…]
As mentioned, the bounds for the distance function apply to obstacles that can be represented by a convex set. In general, a scenario will, however, consist of non-convex obstacles. But it is known that a non-convex set can be represented in general as the union of convex sets [19, 23], and each of them will define a Voronoi region in which the distance function $d(l(t);\Omega)$ is convex. Therefore, the approximation in (11) will be valid except for the segments that cross the boundaries between these regions, the cut locus.
[…]
$\text{error}_{\text{relative}}\approx\frac{\Delta O}{O(s_{i},s_{i+1})}=\frac{L^{2}/2}{\left(\frac{d(s_{i})+d(s_{i+1})}{2}\right)L}=\frac{L}{d(s_{i})+d(s_{i+1})}.$ (12)
Thus, the relative uncertainty of the approximation increases proportionally with L for fixed d(si)+d(si+1), reducing the reliability of the EDF-based term for long segments. This also implies that L is an important parameter, as longer segments not only lead to larger approximation errors but also increase the likelihood of crossing the cut locus. This demonstration is relevant because the value of $L$ depends on the line of sight considered, that is, the distance between $s_{i}$ and $s_{i+1}$ nodes will always be less than or equal to the maximum line of sight considered. So, the shorter the line of sight or length of the segments, the more similar the bounds are. This will be evaluated in Section 5.
Additionally, the connection with the modification introduced in Section 2 regarding limiting the line of sight can also be understood from the behavior of the EDF-based cost term in (2). Since in (2) the EDF-based cost appears in the denominator, the safety-related penalty term is approximately:
$\frac{c_{w}}{O(s_{i},s_{i+1})}\approx\frac{c_{w}}{\left(\frac{d(s_{i})+d(s_{i+1})}{2}\right)L}=\frac{2c_{w}}{(d(s_{i})+d(s_{i+1}))\,L}.$ (13)
Therefore, for fixed values of $d(s_{i})$ and $d(s_{i+1})$, the contribution of the EDF-based cost term scales as $1/L$. This implies that as $L$ increases, the penalty associated with proximity to obstacles decreases inversely with segment length, causing the Euclidean distance to dominate the total cost $c(s_{i},s_{i+1})$ in (2). Consequently, the influence of the EDF term diminishes for long segments, reinforcing the need to limit the line of sight length so that the safety contribution remains meaningful within the cost function.
@@ 4 Selection of visibility neighbours
where $EDF(s)$ and $EDF(s^{\prime})$ are the distances to the closest obstacles in each node $s$ and $s^{\prime}$, and distancedistance is the distance between $s$ and $s^{\prime}$. Notice how positive values indicate that we are getting closer to an obstacle in such direction, while negative ones imply moving away from the obstacle.
Then, the minimum value of $EDF^{\prime}(s^{\prime})$ of the twenty six neighbours, $s_{edf}$, is chosen with the $ComputeMinDerivEDF$ function in Algorithm 1 (see red arrow in Figure 4(a)). On the other hand, the direction of $s$ to the goal is also computed (blue arrow in Figure 4(a)). Once $s_{edf}$ and $s_{goal}$ are computed, the $ComputeVector$ function in line 16 of Algorithm 1 calculates the sum of their corresponding direction vectors and normalizes the result to obtain a unit vector $\hat{u}$ (green arrow in the enlarged part of Figure 4(a)). This vector defines the direction toward the best candidate node among the 26 visibility neighbours to be expanded (green arrow is the direction and dark gray cell the best candidate to expand in Figure 4(b)). Following the computation of $\hat{u}$, $ChooseNeighbours$ function in line 17 in Algorithm 1 selects the neighbours around the best candidate (blue cells in Figure 4(b)). For example, if the FS-Planneris configured to explore 9 neighbours out of the 26 surrounding node $s$, the vector $\hat{u}$ identifies the best candidate to expand, and the remaining 8 neighbours correspond to the cells surrounding this candidate (dar gray cell and blue cells in Figure 4(a). Figure 4 illustrates the selection process when 9 neighbours are chosen for expansion. Finally, the cost, $g$, and parent of each node is computed for the selected neighbours in line 18-23 in Algorithm 1.
@@
Figure 4: Visibility-neighbour selection process. (a) The neighbour with the minimum value of $EDF^{\prime}(s^{\prime})$ is identified ($s_{edf}$) by the ComputeMinDerivEDF function, indicated by the two red arrows, together with the direction to the goal $s_{g}oal$ (blue arrow). These two vectors are combined by the ComputeVector function to obtain the unit direction vector $\hat{u}$ (black arrow). (b) Once the best candidate neighbour (dark gray cell) is selected, the ChooseNeighbours function generates the reduced set of visibility neighbours; in this example, nine neighbours are selected (blue cells).